OpenStudy (buttermequeasy):

Simplify by combining like terms. 6(g^2+5)+2(h^2-3g^2)-8h^2

OpenStudy (anonymous):

answer is 30-6h^2

jhonyy9 (jhonyy9):

6g2 +30 +2h2 -6g2 -8h2 = 30- 6h2 =6(5-h2)